OpenStudy (idealist10):

Consider a set of n antennas of which m are defective and n-m are functional and assume that all of the defectives and all of the functionals are considered indistinguishable. How many linear orderings are there in which no two defectives are consecutive?

OpenStudy (idealist10):

@ganeshie8 @PaxPolaris @sangya21 @Compassionate @uri

OpenStudy (paxpolaris):

freaky

ganeshie8 (ganeshie8):

lol the phrasing is freaky but the problem is simple

ganeshie8 (ganeshie8):

its equivalent to below problem : you have `n` students of which `m` are girls, find the number of possible arrangements such that no two girls are together

OpenStudy (paxpolaris):

if it didn't matter where the defectives go ... the answer would simply be (n choose m)

OpenStudy (paxpolaris):

if m > n/2 ... the answer would be 0

OpenStudy (idealist10):

@ganeshie8

OpenStudy (paxpolaris):

i could recursively calculate that for any number of n and m ... but no idea what the general formula would be :(

OpenStudy (paxpolaris):

*correction if m>(n+1)/2 the answer would be 0. if m=(n+1)/2 the answer would be 1.

OpenStudy (paxpolaris):

\[\text{if }m=2, \to {n\choose 2}-(n-1)={n-1 \choose 2}\]\[\text{if }m=3, \to {n-2 \choose 3}\text{ for }\]\[\text{if }m=4, \to {n-3 \choose 4}\] general formula: ((n-m+1) choose m)

OpenStudy (paxpolaris):

what's the simple way to get this ? @ganeshie8

OpenStudy (paxpolaris):

@perl

OpenStudy (perl):

so basically you have m F's and n-m S's

OpenStudy (perl):

set { F F F F F , S S S S S } and you want no two F's together

OpenStudy (perl):

FSFSFSFSSS , for example

OpenStudy (perl):

lets use smaller numbers, { F F S S SS } F S F S SSS so you have this F S F , that is moving 1 step to the right, how many steps is it moving?

OpenStudy (perl):

do you see how 'FSF' is moving F S F S S S S F S F S S S S S F S F

ganeshie8 (ganeshie8):

this should work \[\rm \binom{n-m+2}{m} (n-m)!\] Notice, the problem makes sense only when \(\rm m >= \frac{n}{2}+1\)

OpenStudy (paxpolaris):

the answer is:\[\large{n-m+1 \choose m}\] it has to always be smaller than (n choose m)

ganeshie8 (ganeshie8):

yes that looks good as the objects are indistinguishable